OpenStudy (anonymous):

Distributive property 10 (x+y-z)

OpenStudy (anonymous):

a(b+c)=ab+ac is the distributive property. so here we have 10(x+y-z)=10x+10y-10z

OpenStudy (anonymous):

Thanks :)